Solve for the unknown
2/3/10 = 1.8/r
Please note the 2 , 3 and 10 are on top of eachother and the 1.8 is on top of the little r.

Answer by stanbon(75887)

Solve for the unknown
2/3/10 = 1.8/r
---------
2/30 = 1.8/r
Invert both sides to get:
r/1.8 = 30/2
r = 1.8*15
r = 27
=========
Cheers,
Stan H.